6x+12y=-6 and 2x+5y=0
Answer by mananth(16946)

6x+12y=-6
/6
x+2y=-1
and 2x+5y=0
1 x + 2 y = -1 .............1
2 x + 5 y = 0 .............2
Eliminate y
multiply (1)by -5
Multiply (2) by 2
-5 x -10 y = 5
4 x + 10 y = 0
Add the two equations
-1 x = 5.00
/ -1
x = -5
plug value of x in (1)
1 x + 2 y = -1
-5 + 2 y = -1
2 y = -1 + 5
2 y = 4
y = 2